Schell Company manufactures automobile floor mats. It currently has two product lines, the Standard and the Deluxe.

Schell has a total of $26,440 in overhead. It currently uses a traditional cost system with overhead applied to the product on the basis of either labor hours or machine hours. Schell has compiled the following information about possible cost drivers and its two product lines:

Schell Company Total Quantity/Amount
Consumed by
Standard Floor Mat Line
Quantity/Amount
Consumed by
Deluxe Floor Mat Line
1,050 labor hours 670 labor hours 380 labor hours
7,120 machine hours 2,910 machine hours 4,210 machine hours

Required:
1. Suppose Schell uses a traditional costing system with direct labor hours as the cost driver. Determine the amount of overhead assigned to each product line.
2. Suppose Schell uses a traditional costing system with machine hours as the cost driver. Determine the amount of overhead assigned to each product line.

Required information

Bunker makes two types of briefcase, fabric and leather. The company is currently using a traditional costing system with labor hours as the cost driver but is considering switching to an activity-based costing system. In preparation for the possible switch, Bunker has identified two activity cost pools: materials handling and setup. Pertinent data follow:

Fabric Case Leather Case
Number of labor hours 14,000 9,000
Number of material moves 540 810
Number of setups 48 72

Total estimated overhead costs are $203,550, of which $168,750 is assigned to the materials handling cost pool and $34,800 is assigned to the setup cost pool.

Required:
1.
Calculate the overhead assigned to the fabric case using the traditional costing system based on direct labor hours.
2. Calculate the overhead assigned to the fabric case using ABC.
3. Was the fabric case over- or undercosted by the traditional cost system compared to ABC?
Answer #1

1a) Overhead assigned to under direct labor hour

Standard Floor Mat Line Deluxe Floor Mat Line
Overhead assigned 26440/1050*670 = 16871 9569

1b) Overhead assigned to under machine hour

Standard Floor Mat Line Deluxe Floor Mat Line
Overhead assigned 26440/7120*2910 = 10806 15634

2a) Overhead assigned to Fabric under traditional costing = 203550/23000*14000 = 123900

2b) Overhead assigned to Fabric under ABC = (540*125+290*48) = 81420

2c) Over costed under Traditional costing by 42480